A box contains 2 bags; bag X and bag Y. Bag X contains 2 red and 1 white beads. Bag Y contains 2 red and 2 white beads. A bag is selected at random. 2 beads are selected, one after the other without replacement. Determine the probability that:
A.2 red beads are selected.
B.1 red and 1 white bead are selected.
C.At least 1 red bead.
1
Expert's answer
2020-08-17T19:10:21-0400

Solution:

Let DXX be the event that the bag X is selected twice.

Let DYY be the event that the bag Y is selected twice.

Let DXY be the event that the bag X is selected once and the bag Y is selected once.


There are 4 options to choose from two bags twice: XX,XY,YX,YY

So,

P(DXX)=1/4

P(DXY)=1/2

P(DYY)=1/4


A. Let A be event that 2 red beads are selected.

Then

P(A)=P(A|DXX)P(DXX)+P(A|DXY)P(DXY)+P(A|DYY)P(DYY)


P(A|DXX)=m/n

m=C22=2!/2!/0!=1

n=C23=3!/2!/1!=3

P(A|DXX)=1/3


P(A|DXY)=m/n

m=C12*C12=2!/1!/1!*2!/1!/1!=2*2=4

n=(C13)(C14)=3!/1!/2!*4!/1!/3!=3*4=12

P(A|DXY)=4/12=1/3


P(A|DYY)=m/n

m=C22=1

n=C24=4!/2!/2!=6

P(A|DYY)=1/6


P(A)=(1/3)*(1/4)+(1/3)*(1/2)+(1/6)*(1/4)

P(A)=7/24

The probability that 2 red beads are selected is equal 7/24


B. Let B be event that 1 red and 1 white bead are selected.

Then

P(B)=P(B|DXX)P(DXX)+P(B|DXY)P(DXY)+P(B|DYY)P(DYY)

P(B|DXX)=m/n

m=C12C11=2*1=2

n=C23=3

P(B|DXX)=2/3


P(B|DXY)=m/n

m=C12*C12 + C11*C12=2*2+1*2=6

n=(C13)(C14)=3*4=12

P(B|DXY)=6/12=1/2


P(B|DYY)=m/n

m=C12 C12=2*2=4

n=C24=4!/2!/2!=6

P(B|DYY)=4/6=2/3


P(B)=(2/3)*(1/4)+(1/2)*(1/2)+(2/3)*(1/4)=7/12

The probability that 1 red and 1 white bead are selected is equal 7/12.


C. Let C be event that at least 1 red bead.

Then C=A+B, since A"\bigcap"B="\varnothing"

So

P(C)=P(A)+P(B)

P(C)=7/24+7/12=21/24=7/8

P(C)=7/8

The probability that at least 1 red bead is equal 7/8.


Here Cmn=n!/m!/(n-m)! - binomial coefficient.